# Security Code Review by Large Language Models

### v2

@@ 3.3. Research Procedure of RQ1
3.3.1. LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December, 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through API. For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs. Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings. The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better, thus were employed in our study.
Table 1. Versions and hyper-parameters of 6 LLMs studied
@@ 3.3.2. Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts. The detailed construction strategies of the five prompts are provided in (Anonymous, 2024).
@@
    Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We directly instruct the LLM to review the provided code for security defects and output their description, code line numbers, and fixing solution. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
@@
    Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information. $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information. The LLM is asked to act as a code reviewer for the project of the provided code in the prompt.
@@
    Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction. $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects, without providing specific CWE details.
@@
    Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
3.3.3. Evaluation Metrics
Aiming to measure the performance of LLMs, we adopted the evaluation methodology proposed by Mahajan et al. (Mahajan et al., 2020) to rate the responses that LLMs generate with some adjustments, as detailed in Table 2. First, we excluded the ‘Unavailable’ category because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers, while a lower M-Score signifies less misleading results produced during security defect detection. Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness, and calculated the average of the three performance scores as the final result.
Table 2. Adjusted IHMU-category for LLM responses
@@ 3.3.3. Evaluation Metrics
Helpful
The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
@@
Uncertain
The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
@@ 3.3.4. Baselines
We attempted different tools and learning techniques for detecting security defects to select appropriate baselines for our work. To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like python, but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. In order to obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
@@ 3.3.5. Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. To rate the responses, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then the first and third authors proceeded to label all the rest responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset and found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
@@ 3.3.6. Consistency Calculation
To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, for a given code file under a specific prompt template, an LLM generated three responses. We then calculated the entropy for the four response categories in the three responses and determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
@@ 3.4. Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze quality problems present in these responses. We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (Kabir et al., 2024) and problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. Notably, we excluded the Consistency theme defined in Kabir et al. (Kabir et al., 2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, 10 responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, problem analysis is a multi-level and multi-label classification task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the remaining data analysis was completed by the two authors and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each problem type, we computed its average frequency in three repetitive experiments as the final occurrence rate of such problem in GPT-4 responses.
@@ 3.5. Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is outlined in the following subsections.
@@ 3.5.1. Attributes Construction
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments and obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, along with their factor values, ready for model fitting.
Factor Source Definition Rationale
Token
Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file. Computed using the TikToken (tik, 2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
FileType
Created Type of the code file, i.e., Source or Auxiliary. Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp). The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
SecurityDefectType
Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. (Yu et al., 2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit
Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity
Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976). Computed using the Lizard (liz, 2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
Community
Created Name of the community source of the code file, i.e., OpenStack or Qt. OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components. The LLMs’ performance may differ on data derived from these two communities.
AuthorExperience
Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
AnnotationRatio
Created The ratio of manual annotations in the code file. If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
AnnotationHasCode
Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
AnnotationisSecurityRelated
Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects). Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.

### v3

@@ 3.3. Research Procedure of RQ1
Table 1. Versions and hyper-parameters of 6 LLMs studied.
@@
3.3.1. LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API. For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs. Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings. The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better and thus were employed in our study.
@@ 3.3.2. Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts. Fig. 2 illustrates the construction strategies for the five prompts.
@@
    Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
@@
    Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information. $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information. The LLM is asked to act as a code reviewer for the project using the provided code in the prompt.
@@
    Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction. $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects without providing specific CWE details.
@@
    Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
Figure 2. Construction templates for the five prompts
Table 2. Adjusted IHMU-category for LLM responses.
@@
Helpful (H) The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
@@
Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
3.3.3. Evaluation Metrics
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al. (Mahajan et al., 2020) for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al. because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories detailed in Table 2. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies less misleading results produced during security defect detection. Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness. We calculated the average of the three performance scores as the final result.
@@ 3.3.4. Baselines
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.
To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
@@ 3.3.5. Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset. We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
@@ 3.3.6. Consistency Calculation
To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
@@ 3.4. Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses. We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (Kabir et al., 2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. In particular, we excluded the Consistency theme defined in Kabir et al. (Kabir et al., 2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, ten responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the remaining data analysis was completed by the two authors and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
@@ 3.5. Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
Factor Source* Definition Rationale
Token Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file. Computed using the TikToken (tik, 2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
FileType Created Type of the code file, i.e., Source or Auxiliary. Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp). The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
SecurityDefectType Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. (Yu et al., 2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976). Computed using the Lizard (liz, 2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
Community Created Name of the community source of the code file, i.e., OpenStack or Qt. OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components. The LLMs’ performance may differ on data derived from these two communities.
AuthorExperience Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
AnnotationRatio Created The ratio of manual annotations in the code file. If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
AnnotationHasCode Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
AnnotationisSecurityRelated Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects). Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
    The sources of factors correspond to the descriptions in bold in Section 3.5.1.
@@ 3.5.1. Attributes Construction
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments. We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment and their factor values ready for model fitting.